THE LAMBDA-TYPE DOUBLING TRANSITIONS OF THE CH RADICAL AND ITS ISOTOPIC SPECIES

\underline{SHAZRENE MOHAMED}, MICHAEL C. McCARTHY, PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, U.S.A. and Division of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138, U.S.A.; JOHN M. BROWN, Physical and Theoretical Chemistry Laboratory, South Parks Road, Oxford OX1 3QZ, England.

The lowest lambda-type doubling transitions of the CH radical and its carbon-13 isotopic species have been directly detected at centimeter wavelengths for the first time. Using a Fourier transform microwave spectrometer and a pulsed supersonic molecular beam, the hyperfine-split components in the X^{2}Π_{1/2} state have been measured to high precision (1 kHz) between 3 and 15 GHz. Owing to the high spectral resolution of this technique, the accuracies of the hydrogen and carbon-13 coupling constants have been significantly improved. A full account of our measurements, including a comparison of the production of CH with different precursor gases and attempts to measure the analogous transitions of the deuterated isotopic species (CD), will be presented.